VISTA IOT #6 MEETING 2009-03-26 ================================ Minutes by: M. Rejkuba ====================== March 26, 2009 09:00 Chile, 14:00 Garching -------------------------------------------- ! ! 2009-03-31 minor modifications TSz ! 2009-04-01 minor modifications MAR ! CONTENTS: 0: Presence at the meeting 1: VISTA schedule, current status and planned activities 2: SADT: status and delivery time line 3: Round table (the state of preparation for VISTA in different groups) 4: Summary of Action Items 5: Dates of the next meetings =========== === 0. PRESENCE ======================================================== =========== Paranal - T. Szeifert ................................... [TSZ] Santiago - V. Ivanov ................................... [VIV] Garching - M. Arnaboldi ................................... [MAR] S. Castro ....................................[SCA] W. Hummel ................................... [WHU] D. Macchetto ................................... [DMA] M. Rejkuba ................................... [MRE] via Phone - J. Emerson ................................... [JEM] CASU: J. Lewis, M. Irwin, S. Hodgkin ............. [JLE] ======================================================== === 1. VISTA schedule, current status and planned activities ======== ======================================================== TSZ gave the summary of the schedule of VISTA based on the latest schedule received from VPO only days before the meeting. This schedule is not official yet, but the time lines between handover and start of operations are proposals which are being discussed with Paranal engineering and will be agreed with the ESO management later. Schedule: ----------- - handover PSO to QMUL - second part of May (dates from VPO schedule) - handover ESO - second part of June (dates from VPO schedule) - end of June (June 22 preliminary date): preliminary version of the instrument package (IP), delivery of all manuals, p2pp, SADT, OT - Science Verification (SV) - beginning of July - mid July - freeze of the IP and delivery to Garching for the Phase 2 - Start of operations - October 1 (dates from VPO schedule) In September some time could be reserved for the mirror coating. This would then happen between the phase 2 deadline and the start of operations. Dry runs with possibly some early survey OBs could be done before/after the mirror coating in September. JEM pointed out that the mirror re-coating was not in the VPO schedule, but was requested by ESO. TSZ replied that in September the coating of the VISTA mirror will anyway be 15 months old and due for the regular re-coating. Additionally this is necessary so that ESO staff becomes familiar with the procedures necessary for the future maintenance of the telescope. The above schedule should not be affected by the new problem that is related to the cryogenic unit. This will however require a warm-up of VIRCAM. MAR asked about the status of the mirror positioning and how well the best position can be recovered. JEM: During the intervention all the surfaces were cleaned and the instrument is now re-attached and according to Gavin the mirror is in the same position as before the intervention. Now the positioning and the alignment has to be improved, but this is expected to be done without problems, because all the mechanical obstacles were removed. They were preparing to move the mirror last night. After that they will go back through the to-do punch/action list until mid May. About 3 weeks are estimated to be needed to reach the technical specs and then some time to analyse the data. TSZ reminded that the warm-up and the work on compressors to correct the problem with cooling will require at least two weeks and this is not clear when it will happen. DMA asked about the current status and current work being done. JEM replied that at the moment focus is being verified and all the actions are taken aiming to get the system running again. The question by DMA whether any work has been done on improving the overheads, JEM answered: NO. ===================================== == 2. SADT: status and delivery timeline ========== ===================================== JEM gave the overview of SADT status and expected delivery. Problem with the tiles around the pole positions is not solved yet, but all the rest of the major changes are in there. At the end of April the person working on the code is leaving (end of contract). JEM will ask Thomas Bierwirth (and ESO) to take over further development and maintenance of the tool. A possibility would be that either the developer comes to ESO or Thomas Bierwirth sends someone to Edinburgh. MAR asked about the delivery time-scale of SADT and whether there are any fixed/agreed dates between VPO and ESO for the SADT PAE at ESO garching. JEM said it would be at the end of April. He will call F. Comeron beginning of the next week. At the moment the version is expected to be able to support the VST phase 2, but this needs to be checked by the VST people. In order to fix one other problem and add the VST module, the XML schema had to be changed. This was approved by Thomas Bierwirth after the last VISTA IOT meeting in February. However, now this means that SADT does not talk to the current version of P2PP. TSZ remarked that he thought that only a "irrelevant parameter" (for P2PP) would be added in this change of XML schema. JEM replied that this was true, but now after VST part was added there are more changes. VIV asked if the problem with rotation was solved. JEM: It was solved but was not tested on sky, and cannot be tested on sky before the new P2PP version is released, with the updated SADT schema. There is a possibility that a sign might be wrong. MRE asked if the XML schema was sent to T. Bierwirth. The answer was "not yet". Also the contact with T. Bierwirth to coordinate the meeting with the developer has to be scheduled ASAP. JEM to get in contact with T. Bierwirth and F. Comeron --> AI. =========================== === 3. Round table activities =============== =========================== 3.1 SciOps: TSZ - discussions with Thomas Bierwirth about OT ranking algorithm; OT demo tomorrow should have close to final version of the tool. P2PP and OT are well on track with the schedule. No activities with the telescope. Still to be checked if pipeline is still running and resume sending data once VISTA starts taking them again. TSZ working on the document to formalize the procedure for feedback from/to CASU. This is not part of the QC, it is only to report in cases where some strange data are found. This will be done through the mailing list. Will be formalized. --> AI VIV - nothing new; given that VISTA was in hands of VPO and there was no new schedule TSZ and VIV were scheduled on VLT. The importance of having the schedule of VISTA activities well in advance was re-iterated due to the need to plan the turnos on Paranal. 3.2 VPO: JEM asked if it would be possible to use the output of the Paranal pipeline for quick looks and checks during the commissioning. Is the current version reducing only calibrations or also science data? SCA replied that the Paranal pipeline reduces all the calibrations and only the first pawprint of science OBs. It can be modified, but the rules have to be changed manually on the pipeline machine (DO stopped, etc... and pipeline re-started). There was a short discussion regarding the time loses due to need to go to the telescope to reset boxes, while the observers are sitting in the Control Building. This happened in January during the visit of DMA several times. However, this was most probably due to problem with the cooling system of the camera. This problem has to be fixed. JEM will ask to be given the timescale. Such frequent trips to the telescope to solve problems should not be part of routine operations. 3.3 CASU: JLE: Still playing with the data. Have 2 months worth of data from Dec/Jan commissioning, which will be put online very soon. At the moment there are not many changes for the Paranal pipeline, and some upgrades on the CASU pipeline. TSZ: brought up the issue of sky subtraction: does anything have to be changed in the observing strategy to achieve a good sky subtraction. Is it enough to go with concatenations? How frequently one has to sample the sky? Do we need a new VISTA template for sky acquisition? When is the first draft based on commissioning January data to be expected with the reports on astrometry, calibrations, accuracy of sky subtraction, etc.? Currently there is no report or written document, but only some presentations. JEM: clarified that this part is delivery from VPO. VPO takes the input from CASU. TSZ: These materials (plots and numbers) are necessary for the manuals. Either we get them now (soon) or after the handover we have to take the data and then spend some weeks figuring out the numbers for the manuals. There will be other activities like the finalization of the IP to be done at the same time. It is much better to avoid work pile-up and potential delays by having this now. JLE: said that will write some report - put together plots and short explanations and send to TSZ --> AI Back to the discussion about the sky subtraction: JEM pointed out that having sky observations in the same OB was not possible due to rule that two telescope presets cannot be done in the same OB. How is this done with HawkI? TSZ and VIV gave details: With Hawk-I there are fixed offsets up to few degrees. When offsets are done with VLT in case of small (jitter) offsets the same guide star is kept, in case of large offsets the new guide star is acquired. In some observations guiding and AO is switched off. For VISTA similar could be done. MAR points out that it is not really important how the sky is acquired. The problem is that in the current VISTA pipeline, the sky has to be in the same OB to be recorgnized by the pipeline. If the sky is taken in a separate OBs, then it is not possible to tell the pipeline about it. VIV suggests use of the containers, but TSZ says that the execution times would be too long, and possibly sky exposure would not be useful (not valid any more) - again this depends on variability. JEM asked to have a copy of some HAWK-I OBs and send it to S. Beard to check if something similar can be implemented for VISTA. MRE asked if it would be possible to use for the sky either other detectors or exposures with the same detector but taken at the offset position of another pawprint within the tile. Also question about the maximum size of offset with VISTA. JLE clarifies that the twilight flats have a different color for the background in the science exposure, therefore they do not correct the scinece frame very well. A dark sky is required. The background information from a different cannot be used for the sky subtraction because different detectors have different gains. TSZ: for some bulge fields, and LMC core taking a sky at some offsets is not good enough. Much larger offsets are needed and in this case they are ill defined. A large offset in ra and then in dec, or viceversa, makes the telescope point in different postions. There is urgent need for information on sky stability: how fast is the sky changing and how frequently it has to be observed. We either need this input from CASU or take time to figure this out. JLE: repeats that the report will be sent -> AI. 3.4 QC (WHU): There are not many news since the last IOT in February. Detailed discussions on QC web pages with VIV and some work on documentation. QC plots of individual detectors were requested by Paranal. This will be implemented in cgi scripts that will be generated on the fly. 3.5 EST (MAR): Together with MRE (USD) work has been done on the document describing the user requirements for visualisation tool for survey monitoring. This tool will support both phase 2 and phase 3. Document has been circulated and the next meeting is scheduled for April 3. For the first phase 2 a quick solution (temporary) has been identified based on the currently available tools with small modifications, while the final tool will be developed later, for phase 3. MAR asked CASU and JEM if EST can be put in the mailing list to get info and access to data when they are posted on the web. JLE: OK JEM: will discuss who will get password. MAR asked VIV about the status of the simulations. The last set of simulations showed that the first three surveys are almost fully completed, but the last two have much lower completeness. TSZ pointed out that the simulation is done with hierarchical scheduling algorithms and that this project is getting too expensive and time consuming. MRE asked if the code is well documented and if it would be possible to pass it to someone else or to add another person to help with the necessary modifications. It would be important to re-run the simulation with the currently planned (realistic) scheduling algorithm. MAR added that this should be done before phase 2 in order to provide feedbacks to the lowest ranked surveys projects so that they can improve their completeness rate, VIV: The report would be needed by about beginning of July to have time to discuss it and send to PIs before the phase 2 at the end of July. He will then deliver the simulation report and the code with comments so that this could be taken over by someone else. MAR - activities related to back end and catalogues - SV waiting for the timeline; plans to meet after Easter and start planning TSZ reminded that JEM and W. Sutherland should be kept in the loop for SV. 3.6 Pipeline (SCA): - pipeline will be upgraded to 0.9 this week and this implies new DFS installation WHU asked if the pipeline processed data on the web will be pipeline protected and who will get access? JLE: says that they will be password protected. JEM: does not want the wide distribution of these data. they are commissioning data and should be used by people that need them. MAR: But before there was a question whether EST will get access?? JEM: OK, but only for "EST to work on, this is not UK deliverable". TSZ: Lot of time can be spent in getting the data basis right that large number of astronomers can understand them. This will be critical for June/July. 3.6 USD (MRE): - activities related to OT/P2PP and Requirements for visualisation tool for monitoring has already been covered Question about the time-scale for VST: TSZ: M2 installed, M1 being shipped from Livorno to Antofagasta (DMA). Therefore September looks ok for commissioning. Instrument scientist on Paranal is F. Selman. DMA: By the handover time (mid May) it is important to inform the PIs giving them information what is happening over summer (activities, overheads, SV, ...) and when they might be expected to provide phase 2 material. They will have to organize the work in advance. MRE agrees, but this message will have to be sent, only when we know that the handover can be passed, not before. TSZ reminds that there are two uncertainties in the current schedule - the time needed to warm-up and fix cryogenics and the confirmation from the engineering on the current plan. JEM would like to have information to PIs even to say "until when there will be no start of operations". This should not happen (TSZ, DMA, MRE). DMA stresses that the message should go out with dates, but if someone calls and asks it is ok to say "not before a certain date". DMA informs the IOT that the ESO DG has approved a meeting with the VST PIs for a discussion of the VST public surveys for the Spetember week 7-11. ====================== === 4. Action Items ======================== ====================== Closed AIs: ----------- - Inform TSZ and VIV about the schedule of VPO staffing and activities over the next few months (JEM) - done; reported at the IOT Pending AIs from previous meetings: ----------------------------------- - Monitor execution times & survey completion status (WIP: MAR,MRE) - Iterate draft of implementation plan (TSZ) - Formalization of the procedure of feedback from CASU about the image quality of science data (TSZ) - Implement SADT change request from PIs about rotation of focal plane into SADT and P2PP (JEM together with T. Bierwirth and P. Nunes) - this is partly done, but T. Bierwirth and P. Nunes have to be contacted still and XML schema sent New action items: ----------------- - JEM: get in contact with T. Bierwirth and P. Nunez for exchange of knowledge with the SADT developer before he leaves at the end of April - JEM: get in contact with F. Comeron to formalize the SADT delivery. - JLE: send CASU report to TSZ; particularly critical sky subtraction issue, but also needed information about the astrometry, calibration accuracy, and other items that should be covered in the User Manual - VIV: by the end of June send the new simulation report and code to be taken over for future simulation - MAR/MRE: contact PIs at the end of May (once the handover is done/certain) and inform them about the timeline and activities. Contact PIs for VST meeting in September - TSZ/VIV: send to JEM information on how Hawk-I deals with sky offsets: templates/OBs? =========================== === 5. Next meeting =============== =========================== Dates proposed by TSZ: ----------------------- 28 Apr. 2009 NN 15:00 - 17:00 (09:00 - 11:00 Chile) 26 Mai 2009 NN 15:00 - 17:00 (09:00 - 11:00 Chile) 30 June 2009 NN 15:00 - 17:00 (09:00 - 11:00 Chile) 28 July 2009 NN 15:00 - 17:00 (09:00 - 11:00 Chile) 25 Aug. 2009 NN 15:00 - 17:00 (09:00 - 11:00 Chile) 22 Sep. 2009 NN 15:00 - 17:00 (09:00 - 11:00 Chile) Please feel free to propose special topics for the agenda.